OpenStudy (anonymous):

Marissa has a snow cone cart that serves about 30 people each hour at the local beach. If the cone shaped snow cone cups have a height of 3 inches and a radius of 1.5 inches, what volume of shaved ice is needed each hour? Round to the nearest cubic inch. Use 3.14 for \pi

Mehek (mehek14):

\(πr^2\dfrac{h}{3}\)

Mehek (mehek14):

\(1.5^2=?\)

OpenStudy (anonymous):

@Aliypop

OpenStudy (anonymous):

are u there

OpenStudy (anonymous):

yes

Mehek (mehek14):

then tell me what \(1.5^2=?\)

OpenStudy (anonymous):

2.25

Mehek (mehek14):

\(π*2.25\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\pi=3.14\]

OpenStudy (anonymous):

so the answer is 2.25 ?

OpenStudy (anonymous):

no do 3.14*2.25

Mehek (mehek14):

no

OpenStudy (anonymous):

7.065

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now do 3/3

OpenStudy (anonymous):

i got 1

OpenStudy (anonymous):

yep now do 7.065*1

OpenStudy (anonymous):

7.065

OpenStudy (anonymous):

yes but that is only for one and they need to know for 30 so do 7.065*30

OpenStudy (anonymous):

211.95

OpenStudy (anonymous):

yes now round to the nearest ones place

OpenStudy (anonymous):

sorry 21.195

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

no it was 211.95

OpenStudy (anonymous):

after rounding it is 212

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so the answer is 212

OpenStudy (anonymous):

Don't forget to medal and fan if you already haven't. Plus send me a message if you need more help. @Aliypop